OpenStudy (anonymous):

Find the six trigonometric functions of theta= -270 degrees

OpenStudy (anonymous):

sin -270 = 1 csc -270 = 1/sin -270 = 1 cos -270 = 0 sec -270 = 1/cos -270 = undefined tan -270 = undefined cot -270 = 1/tan -270 = 0

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

How to make it like this, sin theta = 270 over 1

OpenStudy (anonymous):

its just 270

OpenStudy (anonymous):

270/1 is the same as 270

OpenStudy (anonymous):

@monblake look at the first thing I wrote

OpenStudy (anonymous):

Thank you very much. It helps a lot.